OpenStudy (anonymous):

what does it mean when they ask u to find the product of 237 and 14

OpenStudy (anonymous):

Find what 237 and 14 is. Like 237 + 14 = x? Solve for x.

OpenStudy (anonymous):

Ah I meant 237 x 14 = x.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Yes you are correct.

OpenStudy (anonymous):

i havent been to school in a long time so bare with me please..........could u tell me what they mean by average these numbers 65 49 24 15 also what it means to quotient 225 and 15 please

OpenStudy (anonymous):

To average those numbers (in other words) means to find the mean. To find the mean in a set of numbers you have to add up all the numbers, then divide by how many numbers there are. Like you have 4 numbers. Just add them all together and divide by 4. Also (your second question) is the question like; "What is the quotient of 225 and 15?" Well that just means to divide 225 by 15. Like 225/15 or 225÷15.

OpenStudy (anonymous):

\[237*14=3318 \]

OpenStudy (anonymous):

(200+30+7)*(10+4) (230+7)*(10+4)=2300+920+70+28=3318

OpenStudy (anonymous):

i forgot how do u add fractions like 5/16 and 7/8 ??

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Product_(mathematics)

OpenStudy (anonymous):

1/8(5/2+7)=1/8(19/2)=19/16

OpenStudy (anonymous):

>i forgot how do u add fractions like 5/16 and 7/8 ?? You have to the find common denominator of 5/16 and 7/8. (The bottom number) To make them add(able)/subtract-able you have to get a number on the bottom that everyone shares. You're going to multiply 7/8 by 2 (on top and bottom) to get 14/16. Then you would proceed adding them 5/16 + 14/16. You would only add the numerators (top numbers) and get 19/16.

OpenStudy (anonymous):

what if its more than two fractions like 1 2/3, 1/2, 3/4 how would i do that

OpenStudy (anonymous):

You would find a common denominator for all of them. But I would first make them a mixed fraction. First of all that 1 represents a whole. So take the denominator and add it to the top(numerator) and get rid of the one. You would get 5/2 , 1/2, 3/4. Now find a common denominator (which is four), and multiple the numerator and denominator with a number to get it to four. 5 x 2 / 2 x 2 and 1 x 2 / 2 x 2. ( 3/4 is fine the way it is) so you would eventually get : 10/4 + 1/4 + 3/4 Which would be 14/4 or reduced 7/2

OpenStudy (anonymous):

can you help me on 3.2 is 50% of what???

OpenStudy (anonymous):

50% is a half. So the question then becomes "3.2 is half of what?" As the inverse of a half is 2, we would multiply 3.2 by 2 to get 6.4. So 3.2 is 50% of 6.4.